4x - y, = 2x + y. Solve the system of differential equations with initial conditions x(0)=1, y(0)=2.
Answer #1

8. The ginen system of differential equations are de et = 4x - y e dy = 2x+y » (0-4) x + y = 0 ---- (1) - 2x + (0-1)y = 0 ---2x = 9 e 2t + RG est a = a 2t + Geet The general solution of the system of differential equations are cot ni x(t) = q e 2t +